OpenStudy (anonymous):

Solve. Check each answer. 1. 8 + 5x = -2 2. d/4 - 9 = -3

OpenStudy (anonymous):

(Must show work.)

OpenStudy (anonymous):

8+5x=-2 subtract 8 from both sides 5x=-10 divided both sides by 5 x=-2 d/4 - 9 = -3 add 9 to both sides d/4=6 multiply both sides by 4 d=24